OpenStudy (lilai3):

Community Service: Students are involved in a community service. Suppose 1 student is randomly selected to represent the school at a state-wide awards ceremony. Find the probability of each event. Write as a fraction in simplest form. Girls: 15 Boys: 25 6th graders: 20 7th graders: 8 8th graders:12 P(boy) P(not6th grader) P(girl) P(8th grader) P(girl or boy) P(6th or 7th grader) P(7th grader) P(not a 9th grader)

OpenStudy (campbell_st):

40 students p( boy) = 25/40 or 5/8 P(not 6th) = P(7th) + P(8th) P(8th) = 12/40 or 3/10 P(boy or girl) = 1 I'll let you do the rest

OpenStudy (lilai3):

-_- i just needed help on almost all of them..... o.O

OpenStudy (lilai3):

whatevs

OpenStudy (campbell_st):

so lets look st P(not 6th grader) is the same as P(7th Grader) = P(8th Grader) there are 40 students and 8 7th graders P(7th grade) = 8/40 12 8th graders P(8th Grade) = 12/40 then P(not 6th) = 8/40 + 12/40 = 20/40 = 1/2 its that simple

OpenStudy (lilai3):

ok fine.....(wuteposay(.....thank u?